OpenStudy (anonymous):

Which of the following is equivalent to log 25 125? A. 1.951 B. 2 over 5 C. 2 over 3 D. 3 over 2

OpenStudy (anonymous):

I think it's D but on other people and sites are saying it's c , i know logb^y = log y / log b ... so log 125 / log 25 = 1.5 which is 3/2 .. am i right?

OpenStudy (anonymous):

@peterg2001

Nnesha (nnesha):

\[\log_a b = \frac{ \log b }{ \log a }\] change of base formula

Nnesha (nnesha):

yes you are right !!!

Nnesha (nnesha):

nope 3/2 is right people how are saying is 2/3 they are doing opposite change of base formula is log b/log a( base) base should be at the denominator :-(

Nnesha (nnesha):

:-) **

OpenStudy (anonymous):

Thank you so much .. @Nnesha , so it's D right? , because multiple people were saying c for some reason and i just wanted to clarify that ..

Nnesha (nnesha):

yes it's D\[\huge\rm log_{25} 125\] this is your question right so change of base formula \[\frac{ \log 125 }{ \log 25 }\]

OpenStudy (phi):

\[ \log_{25} 125 = x\] if we raise both sides to the base 25, this "undoes" the log on the left side. in other words, another way to write this equation is \[ 125= 25^x \] You might know that 5*5= 25 and 5*5*5= 125 so we can write it as \[ 5^3 = \left( 5^2\right)^x \] or \[ 5^{2x}= 5^3 \] the bases are equal (to 5), so the exponents are equal \[ 2x= 3 \\ x= \frac{3}{2} \] that means we have found \[ \log_{25} 125 = \frac{3}{2} \]

OpenStudy (anonymous):

Thank you so much @phi for your help ,